# Alpaca against Vicuna: Using LLMs to Uncover Memorization of LLMs

### v2

@@ Abstract
In this paper, we introduce a black-box prompt optimization method that uses an attacker LLM agent to uncover higher levels of memorization in a victim agent, compared to what is revealed by prompting the target model with the training data directly, which is the dominant approach of quantifying memorization in LLMs. We use an iterative rejection-sampling optimization process to find instruction-based prompts with two main characteristics: (1) minimal overlap with the training data to avoid presenting the solution directly to the model, and (2) maximal overlap between the victim model’s output and the training data, aiming to induce the victim to spit out training data. We observe that our instruction-based prompts generate outputs with 23.7% higher overlap with training data compared to the baseline prefix-suffix measurements. Our findings show that (1) instruction-tuned models can expose pre-training data as much as their base-models, if not more so, (2) contexts other than the original training data can lead to leakage, and (3) using instructions proposed by other LLMs can open a new avenue of automated attacks that we should further study and explore.
@@ 1 Introduction
Pre-trained Language models are often instruction-tuned for user-facing applications to enable the generation of high-quality responses to task-oriented prompts (Ouyang et al., 2022; Taori et al., 2023; Chowdhery et al., 2023). A significant body of prior work (Carlini et al., 2022; Biderman et al., 2023a; Shi et al., 2023; Mireshghallah et al., 2022) has extensively defined and studied the memorization of pre-training data in base LLMs, raising concerns in terms of privacy, copyright, and fairness. However, there is a limited understanding of how the instruction-tuning process can affect the memorization and discoverability of pre-training data in aligned models. As such, we set out to answer the question Can we use instruction-based prompts to uncover higher levels of memorization in aligned models?’
The current established method of quantifying memorization in LLMs (Carlini et al., 2023) considers a sequence $d$ memorized in a model in a discoverable manner if prompting the model with the original prefix from the pre-training data would yield sequence $d$ (or a sequence similar to $d$, if we are studying approximate memorization; Biderman et al. 2023a). The assumption in the prior work Carlini et al. (2022; 2023) is that using the ground truth pre-training data as context would provide an upper-bound estimate of memorization. Although, there could exist prompts other than the original training data that would elicit higher levels of training data regurgitation.
@@
Figure 1: Overview of our method: we first create an initial prompt that takes the target training sequence we are probing for and turns it into an instruction. The attacker LLM then uses this prompt to propose multiple candidate prompts that would propel the victim LLM to generate a response that overlaps highly with the training data. We then score each proposed candidate prompt based on two objectives: (1) how much overlap the victim response has with the ground truth training data (the memorization measure, higher, better) and (2) how much overlap the prompt has with the training data (we want this overlap to be small so as not to spill the solution in the instruction). We use this score as a feedback signal for the attacker to optimize the prompt and propose multiple new prompts for the next round of optimization.
To find such prompts, we propose a new optimization method, depicted in Figure 1, where we use another aligned language model as an ‘attacker’ which proposes prompts that would induce the victim (target) model to output a generation that is more faithful to the training data. In this setup, the attacker model iteratively refines its proposed prompts to increase the overlap of the victim output with the ground truth. This is inspired by the victim-play line of work in the computer security literature Wang et al. (2023a). To disincentivize the attacker from feeding the solution to the victim model, we add an extra term to the objective, which minimizes the overlap between the proposed prompts and the target training sequence.
To create robust benchmarks for the evaluation of our approach, we draw a parallel between safety jailbreaking techniques and training data extraction. We leverage automatic prompt optimization to discover prompts that guide the model toward generating outputs closely aligned with its training data. We emphasize that this is different from jailbreaking, as our goal is not to bypass a specific safety feature that prevents training data regurgitation behavior from the model. In our evaluation, we scrutinize the Greedy Coordinate Gradient (CGC; Zou et al. 2023), a white-box prompt optimization technique for identifying prompts that induce detrimental model behaviors. Additionally, we compare our proposed methods against Reverse-LM (Pfau et al., 2023) and sequence extraction (prefix-suffix; Carlini et al. 2022; 2021) across both base models and instruction-tuned models, providing insights on how these widely used methods fare in the context of instruction-tuned models.
We run our method and the baselines on Llama-based, OLMo, and Falcon models (Touvron et al., 2023; Penedo et al., 2023; Groeneveld et al., 2024), and their instruction-tuned variations, including Alpaca (Taori et al., 2023), Tulu (Wang et al., 2023b), and Vicuna (Chiang et al., 2023), spanning $3$ different sequence lengths ($200$, $300$ and $500$) and $5$ different pre-training data domains (following methodology of Duan et al. 2024). Our key contributions and findings are summarized as follows:
  • We propose a black-box prompt optimization approach, tailored for instruction-tuned models, that uses an attacker LLM and shows that our approach uncovers $23.7\%$ more memorization of pre-training data in instruction-tuned models, compared to the prior dominant approach of directly prompting the model with original prefixes from the data Carlini et al. (2022).
  • We also compare the discoverable memorization of pre-training data in instruction-tuned LLMs and their base counterparts and show that using the prior prefix-suffix approach instruction-tuned models demonstrate lower memorization, creating a false sense of higher privacy/lower-risks in these models. Our method, on the contrary, uncovers 12.4% higher memorization in instruction-tuned models, showing that contexts other than the original pre-training data can also lead to leakage, and pointing at the need for better alignment, in terms of privacy.
  • Our experimental results demonstrate that our black-box approach uncovers 12.5% more memorization than the white-box method, GCG, in terms of training data reconstruction overlap.
  • We find that leveraging an open-source model as an attacker can often surpass using a robust commercial model by 2.4%.
We hope that our results and analysis encourage future research to further automate the process of auditing and probing models using other LLMs and to propose more principled, efficient approaches for the reconstruction of training data.
@@ 3 Using LLMs to Probe Memorization in other LLMs
In this section, we begin by formally outlining the optimization problem and specifying our objective function. We present our method’s pipeline, see Figure 1 and Algorithm 1, which includes initialization, sampling, and refinement, creating the optimized prompt.
3.1 Formalizing the Optimization Problem
Consider a sequence $d\in D$, where $D$ is the pre-training dataset of a model $M$. The objective is to find an input prompt $p^{*}$ that the overlap between the output sequence of the model $M(p^{*})$ and $d$ is maximized. Formally, the optimization problem can be expressed as:
@@ 3.1 Formalizing the Optimization Problem
Where $\mathcal{O}_{d,M}(p)=LCS(M(p),d_{\text{suffix}})$ is the objective function we want to maximize for a fixed model $M$ and sequence $d$. $M(.)$ is the operation of decoding from the model $M$, conditioned on a given input. $LCS$ is the longest common subsequence that measures the syntactic similarity between sequences we employ ROUGE-L in our case (Lin, 2004).
In practice, however, LLMs have been shown to be able to regurgitate and repeat their inputs (Zhang & Ippolito, 2023; Priyanshu et al., 2023). Therefore, one obvious solution to this problem could be $p=[z||d]$, where $z$ is an instruction like repeat. To avoid this shortcut, we re-write the objective $\mathcal{O}$ as the following to de-incentive such solutions:
$\mathcal{O}=\alpha\cdot LCS(M(p),d_{\text{suffix}})+(1-\alpha)\cdot-LCS(p,d_{\text{suffix}})$
We add the second term to penalize solutions that overlap highly with the sequence $d_{\text{suffix}}$. $\alpha$ is a hyperparameter to control how much we allow $d$ to be used. Its value is determined by achieving a trade-off between a high memorization score and a low overlap with the ground truth (see Appendix A for details). This problem is, in effect, discrete optimization, previously tackled using gradient-based techniques (Jones et al., 2023; Zou et al., 2023). However, ROUGE-L is not differentiable, and we assume black-box access to the target models to advocate a realistic scenario, rendering gradient-based methods inapplicable.
To solve this, Algorithm 1 shows how we sample from the possible distribution of solutions and find the optimal $p^{*}$. In our setting, we use an alternate model $M^{\prime}(.|[instr])$, with a specific instruction $instr$, as an attacker model that proposes prompts $p$. We perform constrained sampling $p_{t}\sim M^{\prime}(.|[instr\|p_{t-1}])$ at time step $t$ from the proposal distribution, where the constraint is to maximize $\text{LCS}(M(p_{t}),d_{\text{suffix}})$, and we do this with rejection sampling (best-of-n) from $M^{\prime}$. In simpler terms, $M^{\prime}$ seeks the optimal prompt to elicit the sequence $d$ or its similarity from the victim model $M$.
@@
5:     $p_{t}\sim M^{\prime}(Instr|p_{t-1},n=24)$ //Sample 24 prompts
@@
9:  $p^{*}=\arg\max({p}_{0},...,{p}_{t})$ //Obtain the highest scoring prompt across the iterations
@@ 3.2 Optimization via Interactive Sampling
Since the instruction-tuned LM is fine-tuned through a question-answering process to match the user’s intentions better, we design our method to use this instruction-following ability and customize our approach to optimize data extraction. As such, we create prompts in the form of instructions or questions that would propel the target model to output the training data points we want to extract, and then we optimize this ‘initial prompt’ iteratively to get better reconstruction.
@@
To create the initial prompt, we need to somehow transform the training data point into a question. This could be done in different ways, however, we leverage LLMs to do this as well. We instruct this LLM with a ‘meta-prompt’, which is: “Given a paragraph snippet, please generate a question that asks for the generation of the paragraph,” along with the pre-training sample. We also add customized instructions to regularize the prompts, such as “Make sure to keep the question abstract” or “Ensure the question is not overly lengthy.” In practice, we use the meta-prompt on GPT-4 to help generate the initial prompt. Still, we show that utilizing other models, such as Mixtral (Jiang et al., 2024), also yields comparable performance (subsection 5.4).
Finally, we assess the alignment between the ground truth and each prompt, prioritizing prompts with minimal overlap compared to our baseline approach. Further explanations on this will follow. Then, we assess how well the answer to the prompts matches the pre-training sample, saving these paired outcomes for later stages of our procedure.
Interactive Loop After receiving the initial prompt, we utilize a two-step strategy to enhance it for optimal output. These steps involve exploration and exploitation: Initially, we generate $k$ prompts from an attacker LM, assess them, and choose the most effective prompt. This procedure is repeated $i$ times, wherein each iteration exploits the best prompt found and then explores new possibilities through k-samples derived from it.
(1) Best-of-n sampling from $M^{\prime}$: During optimization, the meta-prompt text in this stage differs from the one in the initialization stage. Here, we instruct the model to produce an improved rendition of the prompt, specifically with “I have old questions. Write your new question by paraphrasing the old ones” alongside the preceding step’s prompt. Following this, we supply this instruction to the attacker LLM, generating 24 new prompts for each sample and scoring with our objective function. Ultimately, we choose the prompt with the highest score according to our objective function. On top of this prompt, better-quality samples can be created again in the next step.
@@
(2) Refine: To proceed, we designate the improved prompt from the previous iteration as the starting point and repeat the sampling process three times. This aims to produce a refined version of the original prompt, enhancing extraction capabilities and engaging with the attacker LLM using the prompt from the previous iteration. We do constrained sampling $p_{t}\sim M^{\prime}(\cdot\,|\,[\text{instr}\,||\,p_{t-1}])$ at time step $t$, where the constraint is to maximize $\text{LCS}(M(p_{t}),d)$, and we do this with a rejection sampling (best-of-n) from $M^{\prime}$.
@@ 4.1 Attacker & Victim LLMs
Attacker LLMs: Our method relies on harnessing an open-source model Zephyr 7B, an instruction-tuned variant of the Mistral-7B $\beta$ (Tunstall et al., 2023) as the attacker. We also showcase employing more powerful LLMs as attackers in subsection 5.2.
Victim LLMs: We assess the memorization capabilities of instruction-tuned LLMs compared to their base model across various sizes (7B, 13B, 30B) by applying our method on five open-source models of different sizes by employing the instruction-tuned versions of Llama (Alpaca, Tulu, Vicuna) (Touvron et al., 2023), OLMo (Groeneveld et al., 2024), and Falcon (Penedo et al., 2023). By comparing these instruction-tuned models to their base model, we gain insights into the impact of instruction-tuning on memorization. See Appendix D for more details about the models.

### v3

# Using LLMs to Uncover Memorization of LLMs

@@ Abstract
[…]
[Uncaptioned image] [Uncaptioned image]  Alpaca against Vicuna:
Using LLMs to Uncover Memorization of LLMs
Aly M. Kassem1thanks: Equal Contribution  Omar Mahmoud2*  Niloofar Mireshghallah3* Hyunwoo Kim4  Yulia Tsvetkov3  Yejin Choi5  Sherif Saad1  Santu Rana2 1University of Windsor  2A2I2, Deakin University 3University of Washington  4NVIDIA  5Stanford University  kassem6@uwindsor.ca,  o.mahmoud@deakin.edu.au,  niloofar@cs.washington.edu
@@ 1 Introduction
Pre-trained language models are commonly instruction-tuned for user-facing applications to generate high-quality responses to task-oriented prompts (Ouyang et al., 2022; Taori et al., 2023a; Chowdhery et al., 2023). While extensive prior work has investigated memorization in pre-trained base LLMs and its implications for privacy, copyright, and fairness (Carlini et al., 2022; Biderman et al., 2023a; Shi et al., 2023; Mireshghallah et al., 2022), there is limited understanding of how instruction-tuning affects the memorization and discoverability of pre-training data in aligned models. Studies have shown that aligned LLMs can emit training data up to 150× more often than in regular operation (Nasr et al., 2023). To address this gap, we pose the question: Can we use instruction-based prompts to uncover higher levels of memorization in aligned models? The established method of quantifying memorization (Carlini et al., 2023) assumes that a sequence $d$ is memorized if prompting the model with the original prefix from the training data yields sequence $d$ (or a similar sequence for approximate memorization; Biderman et al. 2023a). However, recent findings suggest that prompts other than the original training data may trigger even higher levels of regurgitation (Schwarzschild et al., 2024). To explore this, we propose a new optimization method, illustrated in Figure Figure 1, where an aligned language model acts as an ‘attacker,’ generating prompts that induce a victim (target) model to produce outputs more faithful to the training data. The attacker refines prompts through a feedback loop guided by a reward function that increases the overlap between the victim’s output and the ground truth. This approach is inspired by adversarial methods in computer security literature (Wang et al., 2023a) and has been effective in jailbreaking attacks (Mehrotra et al., 2023a; Zeng et al., 2024; Ramesh et al., 2024).
@@
Figure 1: Overview of our method: we first create an initial prompt that turns the target training sequence into an instruction. The attacker LLM uses this prompt to generate multiple candidate prompts designed to make the victim LLM produce responses that closely match the training data. We score each candidate based on two criteria: (1) the overlap between the victim’s response and the training data (higher is better) and (2) the overlap between the candidate prompt and the training data (lower is better to avoid revealing the solution). This score guides the attacker in optimizing and generating new prompts for further rounds of optimization.
To evaluate our approach, we draw parallels between safety jailbreaking techniques and training data extraction, using automatic prompt optimization to guide models toward outputs aligned with their training data. However, unlike jailbreaking, our goal is not to bypass specific safety features but to examine memorization. We evaluate our method using Greedy Coordinate Gradient (CGC; Zou et al. 2023), a white-box prompt optimization technique, and compare it to methods like Reverse-LM (Pfau et al., 2023) and sequence extraction (prefix-suffix; Carlini et al. 2022, 2021) across both base and instruction-tuned models. Our method was tested on Llama-based, OLMo, and Falcon models (Touvron et al., 2023; Penedo et al., 2023; Groeneveld et al., 2024), and their instruction-tuned variations, such as Alpaca (Taori et al., 2023a), Tulu (Wang et al., 2023b), and Vicuna (Chiang et al., 2023b), using sequences of $200$, $300$, and $500$ tokens from five pre-training data domains (Duan et al., 2024). We find that our approach uncovers $23.7\%$ more memorization in instruction-tuned models compared to the prefix-suffix method (Carlini et al., 2022), which can give a false sense of privacy. Furthermore, our method reveals 12.4% higher memorization in instruction-tuned models, indicating that contexts beyond the original pre-training data can lead to leakage, highlighting the need for improved privacy measures. To demonstrate the real-world applicability of our method, we conduct four case studies: regurgitation of copyrighted material subsection 6.1, privacy auditing of LLMs subsection 6.2, refusal behavior of LLMs subsection 6.3, and the development of a classifier that detects whether a prompt can elicit memorized data without needing access to the response or the memorized content, enabling a more practical attack subsection 6.4. Our method achieved 39% more extraction in copyright-related queries on the Books3, BooksMIA, and NYT datasets (Computer, 2023; Shi et al., 2023; Grynbaum and Mac, 2023), and a 56.6% increase in privacy auditing over the prefix-suffix approach (Eldan and Russinovich, 2023). Additionally, we show that LLMs do not refuse copyright-related queries with our approach, demonstrating high adversarial effectiveness. Lastly, our classifier reliably detects prompts triggering memorized data in our framework without requiring the actual response, proving more practical. We hope these results encourage further research into automated model auditing and probing using LLMs to develop more efficient data reconstruction methods.
@@ 3 Using LLMs to Probe Memorization in other LLMs
In this section, we begin by formally outlining the optimization problem and specifying our objective function. We present our method’s pipeline, as shown in Figure 1 and Algorithm 1, which includes initialization, sampling, and refinement, creating the optimized prompt.
3.1 Problem Formulation
Consider a set of sequences $D=\{d_{1},\dots,d_{N}\}$, where $D$ is the pre-training dataset of the LLM model $M$. A function $f:d\to p^{*}$ is a transformation process that takes a pre-training sequence $d\in D$ and generates an optimized prompt $p^{*}$ that maximizes the overlap between the output sequence of the model $M(p^{*})$ and pre-training sequence $d$:
@@ 3.1 Problem Formulation
where $\mathcal{O}_{d,M}(p)=\text{LCS}(M(p),d_{\text{suffix}})$ is the objective function to maximize for a fixed model $M$ and sequence $d$. $M(\cdot)$ denotes the operation of decoding from the model $M$, conditioned on a given input. LCS is the longest common subsequence that measures the syntactic similarity between sequences, and in our case, we employ ROUGE-L (Lin, 2004).
@@
5:     $p_{t}\sim M^{\prime}(Instr|p_{t-1},n=24)$ //Sample 24
@@
9:  $p^{*}=\arg\max({p}_{0},...,{p}_{t})$ //get the highest over iters
@@
We consider two settings for the proposed attack. The first one is to estimate its empirical upper bound, which is the default assumption throughout the paper. The second one is a practical setting where we don’t use the full sequence either for evaluation or initialization.
1) Empirical Upper-Bound. To better estimate the empirical upper bound of the attack, we assume that we have access to the full sequence $d$, where sequence $d$ is split into $d_{\text{prefix}}$ and $d_{\text{suffix}}$. We use the full sequence for initialization, which will be discussed later, and for feedback in the objective function, which can be directly used to maximize $\text{LCS}(M(p),d_{\text{suffix}})$. However, LLMs have been shown to regurgitate and repeat their inputs (Zhang and Ippolito, 2023; Priyanshu et al., 2023). Therefore, an obvious solution could be $p=[z||d]$, where $z$ is an instruction like "repeat". To avoid this shortcut, we rewrite the objective $\mathcal{O}$ as follows to de-incentivize such solutions:
$\mathcal{O}=\alpha\cdot\text{LCS}(M(p),d_{\text{suffix}})+(1-\alpha)\cdot(-\text{LCS}(p,d_{\text{suffix}}))$
We include the second term to penalize solutions significantly overlapping with the sequence $d_{\text{suffix}}$. The hyperparameter $\alpha$ regulates how much $d$ is utilized, balancing a high memorization score with minimal overlap with the ground truth (see Appendix A for details).
2) Practical Setting.
In practical scenarios where the suffix $d_{\text{suffix}}$ is inaccessible thus, we can not utilize ROUGE-L for feedback. As a result, we use the $d_{\text{prefix}}$ only for prompt initialization and evaluation. We learn a function $C:\mathcal{P}\to\mathcal{L}$, where $C$ is a binary classifier that takes a prompt $p\in\mathcal{P}$ and outputs a label $l\in\mathcal{L}$, where =$\mathcal{L}=\{T,NT\}$ represents the possibility that a prompt would trigger memorized responses or not. Assume we have access to preference data $\mathcal{D}_{\text{pref}}=\{(p,l)\mid p\in\mathcal{P},l\in\mathcal{L}\}$. We will discuss the details of the classifier in subsection 6.4.
@@ 3.2 Optimization via Interactive Sampling
Initialization. To create the initial prompt, the training data point is transformed into a question. We consider two setups where we use the full sequence $d$ or the prefix only $d_{\text{prefix}}$ (see section 7). An initialization function
$I:\{d_{\text{prefix}},(d_{\text{prefix}},d_{\text{suffix}})\}\rightarrow P_{\text{init}}$
is defined, where $\{d_{\text{prefix}},(d_{\text{prefix}},d_{\text{suffix}})\}$ represents either the prefix alone or both the prefix and suffix.
@@
We instruct LLM with a ‘meta-prompt’ along with the pre-training sample. We also add customized instructions to regularize the prompts to keep them abstract and not overly lengthy. We use the meta-prompt on GPT-4 to help generate the initial prompt. Still, we show that utilizing other models, such as Mixtral (Jiang et al., 2024), also yields comparable performance (section 7).
Interactive Loop. Upon receiving the initial prompt, we employ a two-step strategy to optimize it for the best results, involving exploration and exploitation.
In our setting, we use an alternate model $M^{\prime}(.|[\text{instr}])$, with a specific instruction instr, as an attacker model that proposes prompts $p$. We perform constrained sampling $p_{t}\sim M^{\prime}(.|[\text{instr}\|p_{t-1}])$ at time step $t$ from the proposal distribution, where the constraint is to maximize $\text{LCS}(M(p_{t}),d_{\text{suffix}})$. This is achieved with rejection sampling (best-of-n) from $M^{\prime}$.
(1) Best-of-n sampling from $M^{\prime}$: During optimization, the meta-prompt text evolves from its initialization. We instruct the model to paraphrase the previous prompt $p_{t-1}$ and generate a new one. The attacker LLM produces 24 new prompts per sample, which are scored using our objective function (ROUGE if suffix access is available, otherwise via the proposed classifier) as shown in steps 5, 6, and 7 in Algorithm 1. The highest-scoring prompt is selected, ensuring better-quality samples in the next step where we employ refinement.
@@
(2) Refine: To proceed, We designate the improved prompt from the previous iteration as the starting point and repeat the sampling process three times, following step 4 in Algorithm 1. Each iteration incorporates feedback from the victim to refine the prompt, thereby enhancing extraction capabilities and engaging with the attacker LLM using the previous prompt. At time step $t$, we apply constrained sampling $p_{t}\sim M^{\prime}(\cdot,|,[\text{instr},||,p_{t-1}])$, where the constraint is maximizing $\text{LCS}(M(p_{t}),d)$, using rejection sampling (best-of-n) from $M^{\prime}$.
Average Over Three Sequence Lengths (200, 300, 500)
Model Method Github ArXiv CC C4 Books
Mem LCSP Dis Mem LCSP Dis Mem LCSP Dis Mem LCSP Dis Mem LCSP Dis
Alpaca P-S-Inst .270 .124 - .179 .112 - .155 .104 - .143 .114 - .131 .093 -
Reverse-LM .229 .200 .864 .133 .196 .848 .113 .186 .843 .110 .181 .834 .122 .142 .865
Ours .322 .102 .864 .228 .108 .848 .214 .096 .830 .203 .090 .834 .221 .079 .865
Vicuna P-S-Inst .273 .125 - .213 .112 - .205 .114 - .191 .114 - .198 .093 -
Reverse-LM .255 .200 .864 .200 .196 .848 .173 .186 .830 .173 .181 .834 .166 .142 .865
Ours .325 .096 .864 .232 .104 .853 .213 .092 .838 .201 .084 .841 .223 .079 .866
Tulu P-S-Inst .274 .124 - .207 .112 - .170 .106 - .137 .114 - .172 .093 -
Reverse-LM .245 .200 .864 .153 .196 .848 .121 .186 .830 .117 .181 .834 .135 .142 .865
Ours .359 .104 .857 .237 .104 .851 .221 .094 .835 .210 .086 .836 .233 .079 .865
Seq Len Tulu-7B
200 P-S-Inst .298 .125 - .216 .107 - .176 .103 - .140 .111 - .188 .090 -
Reverse-LM .254 .191 .877 .154 .200 .890 .130 .203 .863 .123 .195 .862 .153 .151 .880
Ours .372 .098 .877 .204 .093 .883 .225 .104 .858 .214 .095 .853 .236 .082 .882
300 P-S-Inst .276 .124 - .209 .112 - .174 .106 - .142 .114 - .178 .095 -
Reverse-LM .246 .203 .881 .157 .196 .853 .125 .190 .822 .116 .182 .826 .134 .145 .877
Ours .341 .084 .878 .248 .108 .856 .222 .099 .824 .209 .090 .825 .231 .079 .872
500 P-S-Inst .247 .124 - .195 .117 - .159 .102 - .128 .117 - .149 .095 -
Reverse-LM .233 .204 .833 .147 .192 .803 .107 .164 .805 .112 .167 .814 .118 .129 .838
Ours .363 .129 .814 .260 .112 .809 .216 0.079 .824 .207 .074 .829 .231 0.076 .841
Table 1: Comparison of our method with baselines across pre-training data domains. Mem denotes the memorization score (ROUGE-L), $LCS_{P}$ is input prompt and suffix overlap, and Dis is optimized vs. initial prompt distance. Results are averaged over three sequence lengths on top, and for the Tulu-7B model, we show a breakdown at the bottom. The highest performance within each domain is bolded.
@@ 4.1 Attacker & Victim LLMs
Attacker LLMs: Our method leverages the open-source Zephyr 7B model, an instruction-tuned variant of Mistral-7B $\beta$ (Tunstall et al., 2023), as the attacker due to its exceptional ability to follow instructions and generate text effectively at the time of writing this paper. We also showcase employing more powerful LLMs as attackers (e.g. GPT-4) in section 7.
Victim LLMs: We assess the memorization capabilities of instruction-tuned LLMs compared to their base model across various sizes (7B, 13B, 30B) by applying our method on five open-source models of different sizes by employing the instruction-tuned versions of Llama-1 (Alpaca, Tulu, Vicuna) (Touvron et al., 2023; Taori et al., 2023b; Wang et al., 2023b; Chiang et al., 2023a), , OLMo (Groeneveld et al., 2024), and Falcon (Penedo et al., 2023) since there is a disclosure in their training data. By comparing these instruction-tuned models to their base model, we gain insights into the impact of instruction-tuning on memorization. See Appendix D for more details about the models.
Figure 2: Comparison of our method to the GCG, P-S baseline, and P-S-instruction on the Llama and its instruction-tuned versions. We evaluate different subsets of the pre-training data and observe that our method consistently outperforms the GCG and prefix-suffix baseline.